# Communication and Energy-Aware Multi-UAV Coverage Path Planning for Networked Operations

### v1

@@ Abstract
This paper presents a communication and energy-aware Multi-UAV Coverage Path Planning (mCPP) method for scenarios requiring continuous inter-UAV communication, such as cooperative search and rescue and surveillance missions. Unlike existing mCPP solutions that focus on energy, time, or coverage efficiency, our approach generates coverage paths that require minimal the communication range to maintain inter-UAV connectivity while also optimizing energy consumption. The mCPP problem is formulated as a multi-objective optimization task, aiming to minimize both the communication range requirement and energy consumption. Our approach significantly reduces the communication range needed for maintaining connectivity while ensuring energy efficiency, outperforming state-of-the-art methods. Its effectiveness is validated through simulations on complex and arbitrary shaped regions of interests, including scenarios with no-fly zones. Additionally, real-world experiment demonstrate its high accuracy, achieving 99% consistency between the estimated and actual communication range required during a multi-UAV coverage mission involving three UAVs.
@@ I INTRODUCTION
In the past few years, the deployment of multiple Unmanned Aerial Vehicles (UAVs) has attracted the interest of many enterprise fields, becoming a powerful tool for professionals to acquire data in a fast and efficient way. Multi-UAV systems have revolutionized various industries by enabling complex missions to be carried out quickly and with precisely, making them invaluable in fields such as surveillance [1], photogrammetry [2], agricultural monitoring, precision farming [3], and Search And Rescue (SAR) [4, 5]. By utilizing multiple UAVs simultaneously, entire regions of interest can be surveyed quickly and thoroughly, addressing challenges that were previously difficult to overcome.
@@
Given the diverse applications of multi-UAV systems, efficient and safe path planning is crucial for executing coordinated missions. A major challenge in Multi-UAV Coverage Path Planning (mCPP) is generating paths that maximize coverage within the Region of Interest (ROI), minimize energy consumption, and maintain communication throughout the mission. Optimizing all of these factors is an NP-hard problem. Although various solutions have been proposed, they include algorithms that handle No-Fly Zones (NFZs) [9, 10, 11], enhance energy efficiency [10, 12], optimize flight speed [10], and support uneven load distribution [9, 12].
@@
While research on mCPP has predominantly focused on optimizing energy, coverage, and time, the consideration of communication requirements in coverage path planning remains largely unexplored. In multi-UAV cooperative missions, reliable communication between vehicles is crucial. Designing coverage paths without accounting for communication needs can lead to frequent disconnections during the mission. This can result in coordination challenges and potential data loss. Maintaining reliable communication is especially critical in operations such as security surveillance, inspection, and SAR, where it is essential for mission success.
@@
In this paper, we propose a communication and energy aware mCPP algorithm that generates coverage paths with minimal inter-UAV communication range requirements. This ensures that all UAVs remain connected throughout the mission in a mesh network architecture. Simultaneously, the algorithm minimizes energy consumption, maximizes area coverage, and completes the mission efficiently. The algorithm utilizes the groundwork laid by Divide Areas based on Robots Initial Positions (DARP) algorithm [9], solving the mCPP problem through area descritization with optimal allocation and constructing a Spanning Tree Coverage(STC) paths on polygonal regions. Energy consumption for the coverage paths is estimated using a direct power consumption model based on the UAVs’ physical parameters.
@@
The primary contribution of this paper is an optimization algorithm that generates coverage paths with minimal inter-UAV communication range requirements. We achieve this by treating the range requirement and energy consumption as hyperparameters in the optimization process, solved using Bayesian optimization techniques across multiple DARP initialization trials. Additionally, we optimize the starting points on generated paths to further minimize communication range requirements. A time-sampling methodology is implemented on these paths, followed by continuous connectivity evaluation to analyze range needs. An optimal shift and rotation technique is applied before discretizing the polygon into cells, enhancing coverage efficiency. This multi-layered optimization framework improves inter-UAV connectivity and energy efficiency, while also supporting complex polygons with NFZs.
@@ II RELATED WORK
Figure 1: Optimization process breakdown: ROI Discretization through optimal shifts and rotation, DARP-MST path generation, time-sampling & energy calculation, start point optimization by connectivity analysis, hyper parameter optimization with range and energy for next DARP initialization
@@
Among the key studies, Comprehensive analyses of CPP techniques are provided in [16, 17], where the authors explore a range of bio-inspired, computational, and analytical CPP methodologies. These studies offer valuable insights into area division and path-planning algorithms for both simple and complex ROIs. They address CPP challenges in both 2D and 3D environments, with and without no-fly zones, primarily focusing on single-robot applications while also considering multi-robot path planning. The use of boustrophedon cell division [17], STC [9], and energy-aware CPP strategies [di2015energy] are highlighted, with emphasis on optimizing flight parameters such as energy consumption, path length, and coverage efficiency. These analyses underscore the complexities of CPP, particularly in cooperative multi-robot scenarios.
Another notable contribution is the introduction of boustrophedon cell decomposition as an alternative to classical trapezoidal decomposition [11, 18]. This technique enables CPP in ROIs with no fly zone by partitioning them into cells and using systematic back-and-forth sweeping paths within each cell. A graph traversal method is then used to connect these individual paths into a complete coverage path. Building on this, [19] revisits boustrophedon CPP by replacing the graph traversal method with the Generalized Traveling Salesman Problem (GTSP) [20] to determine the optimal path configuration, minimizing time and directional changes and improving energy efficiency.
@@
Comm. aware Energy aware Energy efficient Custom load dist. No fly zones
@@
All of the discussed CPP solutions utilize area partitioning and optimal allocation. In [25], a formation flying technique was proposed for multi-UAV coverage missions in ROIs without any NFZs. This approach ensures that the inter-UAV distance remains within the range due to the rigid formation shape, thereby maintaining the UAVs within communication range of each other. Another distinct technique involves discretizing the ROI into grid cells and performing STC. In [9], a fast and effective algorithm is proposed by discretizing the ROI into square grid cells and then using a modified Voronoi algorithm to allocate cells to each robot. The STC algorithm [29] is then implemented for each region to generate independent coverage paths. A similar approach is used in [30], where adaptive grid size discretization and the optimal allocation minimum spanning tree are employed to generate smooth and minimal-cost paths.
To optimize energy in DARP-generated paths, [12] proposes the use of machine learning techniques for generating energy-efficient coverage paths. In this approach, the energy consumption of paths is mapped to the number of turns, and optimization techniques are applied to generate paths with fewer turns across different DARP initialization.

### v2

@@ Abstract
This paper presents a communication and energy-aware Multi-UAV Coverage Path Planning (mCPP) method for scenarios requiring continuous inter-UAV communication, such as cooperative search and rescue and surveillance missions. Unlike existing mCPP solutions that focus on energy, time, or coverage efficiency, the proposed method generates coverage paths that require minimal communication range to maintain inter-UAV connectivity while also optimizing energy consumption. The mCPP problem is formulated as a multi-objective optimization task, aiming to minimize both the communication range requirement and energy consumption. Our approach significantly reduces the communication range needed for maintaining connectivity while ensuring energy efficiency, outperforming state-of-the-art methods. Its effectiveness is validated through simulations on complex and arbitrary shaped regions of interests, including scenarios with no-fly zones. Additionally, real-world experiment demonstrate its high accuracy, achieving 99% consistency between the estimated and actual communication range required during a multi-UAV coverage mission involving three UAVs.
@@ I INTRODUCTION
In the past few years, the deployment of multiple Unmanned Aerial Vehicles (UAVs) has attracted the interest of many enterprise fields, becoming a powerful tool for professionals to acquire data in a fast and efficient way. Multi-UAV systems have revolutionized various industries by enabling complex missions to be carried out quickly and with precisely, making them invaluable in fields such as surveillance [1], photogrammetry [2], agricultural monitoring, precision farming [3], and Search And Rescue (SAR) [4, 5]. By utilizing multiple UAVs simultaneously, entire Regions Of Interest (ROI) can be surveyed quickly and thoroughly, addressing challenges that were previously difficult to overcome.
@@
Given the diverse applications of multi-UAV systems, practitioners in the field face significant challenges in maintaining reliable communication networks while executing missions efficiently. In real-world scenarios like coordinated SAR operations, UAVs must continuously share critical information such as victim locations and environmental hazards. These operations are complicated by the limited communication range of onboard modules, which can lead to network fragmentation when UAVs disperse. Traditional star topologies compound this issue by introducing latency and creating a single point of failure, while mesh architectures provide greater reliability through direct peer-to-peer communications and adaptation to dynamic topology changes. These communication requirements directly impact path planning strategies. Consequently, a major challenge in Multi-UAV Coverage Path Planning (mCPP) is generating paths that not only maximize coverage within the ROI and minimize energy consumption, but also maintain communication links throughout the mission. Optimizing all of these factors is an NP-hard problem. Although various solutions have been proposed, they include algorithms that handle No-Fly Zones (NFZs) [9, 10, 11], enhance energy efficiency [10, 12], optimize flight speed [10], and support uneven workload distribution [9, 12].
@@
While research on mCPP has predominantly focused on optimizing energy, coverage, and time, the consideration of communication requirements in coverage path planning remains largely unexplored. In multi-UAV cooperative missions, reliable communication between vehicles is crucial. Designing coverage paths without accounting for communication needs can lead to frequent network disconnections during the mission. This can result in coordination challenges and potential data loss. Maintaining reliable communication is critical in operations such as security surveillance, inspection, and SAR, where it is essential for mission success.
@@
In this paper, we propose a communication and energy aware mCPP algorithm that generates coverage paths with minimal inter-UAV communication range requirements. This ensures that all UAVs remain connected throughout the mission in a mesh network architecture. Simultaneously, the algorithm minimizes energy consumption, maximizes area coverage, and completes the mission efficiently. The algorithm utilizes the groundwork laid by Divide Areas based on Robots initial Positions (DARP) algorithm [9], solving the mCPP problem through area discretization with optimal allocation and constructing a Spanning Tree Coverage(STC) paths on discretized regions. Energy requirement for the coverage paths is estimated using the direct power consumption model from the UAVs.
@@
The primary contribution of this paper is an optimization algorithm that generates coverage paths with minimal inter-UAV communication range requirements. We achieve this by treating the range requirement and energy consumption as hyperparameters in the optimization process, solved using Bayesian optimization techniques across multiple DARP initialization trials. Additionally, we optimize the launch points on generated paths to further minimize communication range requirements. A time-sampling methodology is implemented on these paths, followed by continuous connectivity evaluation to analyze range needs. An optimal shift and rotation technique is applied before discretizing the polygon into cells, enhancing coverage efficiency. This multi-layered optimization framework improves inter-UAV connectivity and energy efficiency, while also supporting complex polygons with NFZs.
@@ II RELATED WORK
Figure 1: Optimization process breakdown: ROI Discretization through optimal shifts and rotation, DARP-MST path generation, time-sampling & energy calculation, launch point optimization by connectivity analysis, Bayesian optimization with range and energy for next DARP initialization
@@
Among the key studies, Comprehensive analyses of CPP techniques are provided in [16, 17], where the authors explore a range of bio-inspired, computational, and analytical CPP methodologies. These studies offer valuable insights into area division and path-planning algorithms for both simple and complex ROIs. They address CPP challenges in both 2D and 3D environments, with and without NFZs, primarily focusing on single-robot applications while also considering multi-robot path planning. The use of trapezoidal cell decomposition [17], STC [9], and energy-aware CPP strategies [10] are highlighted, with emphasis on optimizing flight parameters such as energy consumption, path length, and coverage efficiency. These analyses underscore the complexities of CPP, particularly in cooperative multi-robot scenarios.
Another notable contribution is the introduction of boustrophedon cell decomposition as an alternative to classical trapezoidal decomposition [11, 18]. This technique enables CPP in ROIs with NFZs by partitioning them into cells and using systematic back-and-forth sweeping paths within each cell. A graph traversal method is then used to connect these individual paths into a complete coverage path. Building on this, [19] revisits boustrophedon CPP by replacing the graph traversal method with the Generalized Traveling Salesman Problem (GTSP) [20] to determine the optimal path configuration, minimizing time and directional changes and improving energy efficiency.
@@
Comm. aware Energy aware Custom load dist. No-fly zones
@@
All of the discussed CPP solutions utilize area partitioning and optimal allocation. In [25], a formation flying technique was proposed for multi-UAV coverage missions in ROIs without any NFZs. This approach ensures that the inter-UAV distance remains within the range due to the rigid formation shape, thereby maintaining the UAVs within communication range of each other. Another distinct technique involves discretizing the ROI into grid cells and performing STC. In [9], a fast and effective algorithm is proposed by discretizing the ROI into square grid cells and then using a modified Voronoi algorithm to allocate cells to each robot. The STC algorithm [29] is then implemented for each region to generate independent coverage paths. A similar approach is used in [30], where adaptive grid size discretization and the optimal allocation Minimum Spanning Tree (MST) are employed to generate smooth and minimal-cost paths.
To optimize energy in DARP-generated paths, (O-DARP) [12] proposes the use of machine learning techniques for generating energy-efficient coverage paths. In this approach, the energy consumption of paths is mapped to the number of turns, and optimization techniques are applied to generate paths with fewer turns across different DARP initialization.